Fama’s Llamas has a weighted average cost of capital of 8.2 percent. The company’s cost of equity is 11.8 percent, and its cost of debt is 6.4 percent. The tax rate is 22 percent. What is the company’s debt-equity ratio? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., 32.1616.)
Answer #1

After-tax cost of debt=6.4*(1-tax rate)

=6.4*(1-0.22)=4.992%

Let debt be $x

Equity be $y

Total=(x+y)

WACC=Respective costs*Respective weight

8.2=(x/(x+y)*4.992)+(y/(x+y)*11.8)

8.2*(x+y)=4.992x+11.8y

8.2x+8.2y=4.992x+11.8y

x=(11.8-8.2)y/(8.2-4.992)

=1.1222y

Hence debt to equity ratio=debt/equity

=1.1222(Approx).